OpenStudy (nuccioreggie):

The work of a student to solve the equation 4(2x − 4) = 8 + 2x + 8 is shown below: Step 1: 4(2x − 4) = 8 + 2x + 8 Step 2: 6x − 8 = 16 + 2x Step 3: 6x − 2x = 16 + 8 Step 4: 4x = 24 Step 5: x = 6 In which step did the student first make an error and what is the correct step? Step 2; 8x − 4 = 2(6 + x + 6) Step 2; 8x − 16 = 16 + 2x Step 3; 6x − 2x = 16 − 8 Step 3; 6x + 2x = 16 + 8

OpenStudy (nuccioreggie):

@AlexandervonHumboldt2

OpenStudy (mathmale):

I'd suggest you begin with the original equation, 4(2x − 4) = 8 + 2x + 8, and go through the actual steps necessary to solve it for x. compare your work, each step of the way, with what this other student has done. Please show your work. Ask questions along the way.

OpenStudy (nuccioreggie):

@mathmale i think its A

OpenStudy (whovianchick):

Step 2. Step 1 is 4(2x − 4) = 8 + 2x + 8 Step 2 is distributing the parenthesis and adding like terms. However, the student multiplied the parenthesis wrong. She said 4(2x-4)= 6x-8. Instead of multiplying the terms, they added them. The correct solution would be 8x-16. Therefore, the answer correct answer choice is B: Step 2; 8x − 16 = 16 + 2x

OpenStudy (nuccioreggie):

thank you can you help me with only 9 more questions

OpenStudy (whovianchick):

Remember, when multiplying parenthesis, you need to multiply the number outside the parenthesis to BOTH values inside. Like this: 4(2x-4) 4(2x) + 4(-4) 8x-16

OpenStudy (whovianchick):

And if you have more questions, post them seperately. It's not fair to the people helping if you post them all here. :) If you have 9 more questions, you should have 9 more medals to give, but you can't if they're all in one chat.

OpenStudy (nuccioreggie):

How many solutions does the equation 6a − 3a − 6 = −2 + 3 have? Two None One Infinitely many

OpenStudy (whovianchick):

Please post new questions.